# Refining local-type primordial non-Gaussianity: Sharpened $b_{\phi}$ constraints through bias expansion

### v1

@@ I Introduction
State-of-the-art numerical simulations have revolutionized our ability to explore the complex interplay between primordial physics and LSS (e.g., 2009MNRAS.396…85D, ; 2010JCAP…10..022W, ; 2023ApJ…943…64C, ; 2023ApJ…943..178C, ; 2024JCAP…02..048F, ; 2024A&A…689A..69G, ). These simulations provide an invaluable laboratory for studying the link between galaxy formation and primordial physics, which is crucial for disentangling $f_{\rm NL}^{\rm loc}$ from $b_{\phi}$. Recent works have explored the relationship between $b_{\phi}$ and $b_{1}$ in simulations 2020JCAP…12..013B ; 2020JCAP…12..031B ; 2022JCAP…01..033B ; 2022JCAP…11..013B and found that beyond the simple linear (universality) relation, there is a significant secondary dependence of $b_{\phi}$ on halo formation history properties such as concentration 2023JCAP…01..023L ; 2024PhRvD.109j3530H . Unfortunately, most of these properties are not observable and instead a number of large-scale analyses only measure ‘bias parameters’ associated with a perturbative expansion model (2004astro.ph.12025T, ).
In this work, we demonstrate that incorporating a detailed understanding of tracer bias parameters into the estimation of $b_{\phi}$ can significantly enhance the precision of PNG constraints. This approach not only sheds light on the complex interplay between astrophysical quantities and cosmological perturbation theories, but also sets the stage for more accurate and robust measurements of the primordial fluctuations in the early Universe. In particular, we aim to refine the estimation of $b_{\phi}$ using a Lagrangian bias expansion model. This approach allows us to incorporate the influence of various bias parameters and halo properties on $b_{\phi}$, thereby providing a more accurate and comprehensive understanding of the scale-dependent bias induced by local-type PNG. We utilize the suite of high-resolution N-body simulations, AbacusSummit, together with the ‘Separate Universe’ technique, which can be employed to understand the response of small-scale galaxy physics to long-wavelength fluctuations. Furthermore, we extend our study beyond halos to galaxies by utilizing an advanced halo occupation distribution (HOD) model to approximate the luminous red galaxies (LRGs) population targeted by the Dark Energy Spectroscopic Instrument (DESI) survey. We show that we are able to improve the uncertainty on $b_{\phi}$ and thus $f_{\rm NL}^{\rm loc}$ by more than 80% for all galaxy tracer samples considered in this study.
@@ II.1 AbacusSummit
Here, we employ the base resolution boxes of AbacusSummit, each of which contains 69123 particles in a $2\,h^{-1}{\rm Gpc}$ box, each with a mass of $M_{\rm part}=2.1\times 10^{9}\,h^{-1}{\rm M}_{\odot}$. While the AbacusSummit suite spans a wide range of cosmologies, here we focus on a single realization of the fiducial cosmology (Planck 2018: $\Omega_{b}h^{2}=0.02237$, $\Omega_{c}h^{2}=0.12$, $h=0.6736$, $2.083010^{9}A_{s}=$ 2.0830, $n_{s}=0.9649$, $w_{0}=-1$, $w_{a}=0$), AbacusSummit_c000_ph000 as well as two additional runs with a 2% higher and lower values of $\sigma_{8}$, AbacusSummit_c112_ph000 and AbacusSummit_c113_ph000. In particular, utilize the halo and particle catalogs (2022MNRAS.509..501H, ) as well as initial conditions outputs. For full details on all data products, see Ref. (2021MNRAS.508.4017M, ).
@@ II.2 Concentration definition
The link between halo concentration and accretion history has been studied extensively in the literature (1997ApJ…490..493N, ; 2002ApJ…568…52W, ; 2014MNRAS.441..378L, ; 2016MNRAS.460.1214L, ), and it has been shown that recent merger activity induces dramatic changes in the halo concentration, and that these responses linger over a period of several dynamical times, corresponding to many Gyr (see, e.g., 2020MNRAS.498.4450W, ). Relevant to galaxy surveys is the fact that halo concentration has a bearing on both the halo occupation distribution and the galaxy clustering clustering (e.g., 2001MNRAS.321..559B, ; 2014MNRAS.441..378L, ; 2015ApJ…799..108D, ; 2014MNRAS.441.3359D, ; 2018MNRAS.474.5143M, ), and recent evidence in data suggests that different galaxy samples have different host halo concentrations (2021MNRAS.502.3582Y, ).
@@ II.3 Local environment
The effect of environment on the halo and galaxy clustering has been studied extensively in the literature (2007MNRAS.378..641A, ; 2017A&A…598A.103P, ; 2018MNRAS.476.5442P, ; 2018MNRAS.473.2486S, ). A halo residing in a dense region is expected to contain more galaxies on average than a halo in an underdense region because halos in overdense regions experience more mergers, whereas those in underdense regions accrete mass more smoothly.
To define halo environment, we adopt the same formalism as Ref. (2020MNRAS.493.5506H, ). Specifically, for each halo, we find all neighboring halos beyond its virial radius but within $r_{\rm max}=5\,h^{-1}{\rm Mpc}$ of the halo center. We sum the mass of all neighboring halos, denoted by $M_{\rm env}$, and define the environment parameter, $f_{\rm env}$:
@@ II.4 Separate Universe approach
We employ the separate universe technique to measure the local-type PNG-induced bias, $b_{\phi}$, by invoking the equivalence between the response of galaxy formation to long-wavelength perturbations and the response of galaxy formation to changes in the background cosmology. The separate universe argument states that as long as the physics of galaxy formation acts on much smaller scales compared with the size of the long-wavelength perturbations, these long-wavelength perturbations act as a modified background to the process of galaxy formation (this is known as the ‘peak-background split’ argument; see e.g., 1984ApJ…284L…9K, ; 1986ApJ…304…15B, ). Thus, the formation of galaxies at fixed cosmology in some region of space embedded in a long-wavelength fluctuation is equivalent to the formation of galaxies in a modified cosmology.
To access the quantity $b_{\phi}$, the modified cosmology needs to have a different amplitude of the primordial scalar power spectrum, $A_{s}$, or equivalently, the amplitude of the linear power spectrum on the scale of $8\ {\rm Mpc}/h$, i.e., $\sigma_{8}$. Mathematically, the PNG-induced bias, $b_{\phi}$, is defined as
@@
where $\phi_{L}$ is to be treated as a constant locally and denotes the amplitude of the long-wavelength potential perturbation. Thus, we can evaluate $b_{\phi}$ in a separate universe with a different value of $A_{s}$ ($\sigma_{8}$). In practice, we typically generate simulations with the same initial seed as the fiducial box, but with an input power spectrum file multiplied by $\left[1+\delta\sigma_{8}^{2}\right]$. In observations, making this measurement directly is extremely challenging, but recent works have suggested a possible path forward that directly extracts the PNG signal from the galaxy density field (2023PhRvD.107f1301G, ).
Here, we employ the ‘Linear derivative’ AbacusSummit boxes, base_c112_ph000 and base_c113_ph000, which have the same initial seed as the fiducial simulation base_c000_ph000, but a different value of $\sigma_{8}$: namely, $\sigma_{8}^{\rm high}=1.02\times\sigma_{8}^{\rm fid}$ and $\sigma_{8}^{\rm low}=0.98\times\sigma_{8}^{\rm fid}$, respectively. We estimate $b_{\phi}$ in the separate universe approach as
@@
where $n_{h}^{\rm fid}(z)$, $n_{h}^{\rm high}(z)$ and $n_{h}^{\rm low}(z)$ is the mean number density of halos (galaxies) in the fiducial, high- and low-$\sigma_{8}$ simulations at some redshift $z$.
@@ II.5 Hybrid Effective Field Theory
In this section, we summarize our approach for estimating the higher-order bias parameters, used to describe the connection between biased tracers such as galaxies and halos, and the underlying matter field in Lagrangian Perturbation Theory (LPT). For a review of perturbation theory, see Ref. 2002PhR…367….1B ; 2018PhR…733….1D ; 2020moco.book…..D .
In the Lagrangian picture, we work with infinitesimal fluid elements labeled by their initial (Lagrangian) positions $\bm{q}$. Their dynamics are encoded in a displacement vector $\bm{\Psi}(\bm{q},\eta)$, sourced by the gravitational potential and defined such that the Eulerian (comoving) positions $\bm{x}$ of the fluid element at some conformal time $\eta$ is $\bm{x}(\bm{q},\eta)=\bm{q}+\bm{\Psi}(\bm{q},\eta)$ 2015JCAP…09..014V ; 2014JCAP…05..022P . The 3D distribution of cosmological observables such as the galaxy field is, in general, determined by complex small-scale astrophysical processes and their response to the large-scale matter distribution (2018PhR…733….1D, ).
We can thus write the dependence of the galaxy field along its trajectory as an expansion to second order in the initial conditions adopting the Lagrangian framework 2009JCAP…08..020M ; 2015JCAP…11..007S ; 2014JCAP…08..056A ; 2016JCAP…12..007V :
@@
where $b_{1}$, $b_{2}$, $b_{s}$ and $b_{\nabla}$ are free bias parameters and $s_{L}^{2}=s_{ij}s^{ij}$ is the shear/tidal field, with $s_{ij}\equiv(\partial_{i}\partial_{i}/\partial^{2}-\delta_{ij}/3)\ \delta_{L}$. Note that the functional $F(\bm{q})$ can only depend on scalar combinations of the density field and the tidal field. Nonlocality effects are handled by performing a Taylor expansion of the $\delta_{L}$ field around $\bm{q}$. To lowest order, this calls for the inclusion of $\nabla^{2}\delta_{L}$. We note that on nonlocal scales, smaller than the scale of astrophysical processes, the Lagrangian approximation breaks down. The functional can then be advected to the real-space (Eulerian) position $\bm{x}$ 2008PhRvD..77f3530M :
@@
Thus, the 21cm clustering signal contains a dynamics piece, $\bm{\Psi}$, as well as a piece depending on the initial conditions, $F(\bm{q})$.
To estimate the bias parameters, $b_{i}$, of our model, we minimize the difference between the model and the observable, $\delta_{\rm g}$. Our result can be expressed through the stochastic field $\epsilon(\mathbf{k})$, which represents residuals after removing deterministic contributions from the model:
@@
We obtain the bias parameters by solving the least squares problem of minimizing the error power spectrum, which yields the following expression for the bias parameters:
@@
Here, we estimate the bias parameters up to some maximum wavenumber, $k_{\rm max}$, as follows:
@@ II.6 Halo Occupation Distribution
The vanilla HOD model for both LRGs and QSOs is given by the standard formalism of Ref. (2005ApJ…633..791Z, ):
@@
where we note that the LRG satellite occupations equation has a modification in the form of multiplication by $\bar{n}_{\mathrm{cent}}(M)$. $M_{\mathrm{cut}}$ determines the minimum mass of a halo to host a central galaxy, $M_{1}$ sets the pivot scale of the power law of satellite occupation, $\sigma$ controls the steepness of the transition from 0 to 1 in the number of central galaxies, $\alpha$ is the power law index on the number of satellite galaxies, $\kappa M_{\mathrm{cut}}$ gives the minimum halo mass to host a satellite galaxy, $f_{\mathrm{ic}}$, which is a downsampling factor controlling the overall number density of the mock galaxies.
In the vanilla model, the velocity of the central galaxy is taken as the average velocity of the so-called “L2” subhalo (see 2022MNRAS.509..501H, ). For the satellite galaxies, the velocities are inherited from random halo particles. The analysis of Ref. (2023arXiv230606314Y, ) also includes velocity bias, which is necessary for modeling redshift-space clustering on small scales (e.g. 2015MNRAS.446..578G, ; 2022MNRAS.510.3301Y, ). In the AbacusHOD model, it is parametrized as:
  • $\alpha_{\mathrm{vel,c}}$ modulates the peculiar velocity of the central galaxy relative to the halo center along the line-of-sight (LOS):
    $v_{\mathrm{cent,z}}=v_{\mathrm{L2,z}}+\alpha_{\mathrm{vel,c}}\delta v(\sigma_{\mathrm{LoS}}),$ (18)
    where $v_{\mathrm{L2,z}}$ denotes the LOS component of the central subhalo velocity, $\delta v(\sigma_{\mathrm{LoS}})$ denotes the Gaussian scatter, and $\alpha_{\mathrm{vel,c}}$ is the central velocity bias parameter.
  • $\alpha_{\mathrm{vel,s}}$ modulates how the satellite galaxy peculiar velocity deviates from that of its host particle:
    $v_{\mathrm{sat,z}}=v_{\mathrm{L2,z}}+\alpha_{\mathrm{vel,s}}(v_{\mathrm{p,z}}-v_{\mathrm{L2,z}}),$ (19)
    where $v_{\mathrm{p,z}}$ denotes the line-of-sight component of particle velocity, and $\alpha_{\mathrm{vel,s}}$ is the satellite velocity bias parameter.

### v2

@@ I Introduction
State-of-the-art numerical simulations have revolutionized our ability to explore the complex interplay between primordial physics and LSS (e.g., 2009MNRAS.396…85D, ; 2010JCAP…10..022W, ; 2023ApJ…943…64C, ; 2023ApJ…943..178C, ; 2024JCAP…02..048F, ; 2024A&A…689A..69G, ). These simulations provide an invaluable laboratory for studying the link between galaxy formation and primordial physics, which is crucial for disentangling $f_{\rm NL}^{\rm loc}$ from $b_{\phi}$. Recent works have explored the relationship between $b_{\phi}$ and $b_{1}$ in simulations 2020JCAP…12..013B ; 2020JCAP…12..031B ; 2022JCAP…01..033B ; 2022JCAP…11..013B and found that beyond the simple linear (universality) relation, there is a significant secondary dependence of $b_{\phi}$ on halo formation history properties such as concentration 2023JCAP…01..023L ; 2024PhRvD.109j3530H . Unfortunately, most of these properties are not observable, and instead, a number of large-scale analyses only measure ‘bias parameters’ associated with a perturbative expansion model (2004astro.ph.12025T, ). Alternatively, machine-learning approaches trained on hydrodynamical simulations have shown to be able to capture some of the dependence of $b_{\phi}$ on secondary parameters and thus improve the constraints on $f_{\rm NL}^{\rm loc}$, while at the same time allowing for a more optimal sample splits for multi-tracer analyses 2023JCAP…08..004S .
In this work, we demonstrate that incorporating a detailed understanding of tracer bias parameters into the estimation of $b_{\phi}$ can significantly enhance the precision of PNG constraints. A similar point was discussed recently in 2024arXiv241018039S , where the authors perform field-level fits of $b_{\phi}$ to perturbative mocks, highlighting the need for a reliable $b_{\phi}$ prior, due to the large residual degeneracy between $b_{\phi}$ and $f_{\rm NL}^{\rm loc}$ even at the field level.
Here, we aim to refine the estimation of $b_{\phi}$ using a Lagrangian bias expansion model to fully non-linear N-body simulations and realistic galaxy samples. This approach allows us to incorporate the influence of various bias parameters and halo properties on $b_{\phi}$, thereby providing a more accurate and comprehensive understanding of the scale-dependent bias induced by local-type PNG. We utilize a suite of high-resolution N-body simulations, AbacusSummit, together with the ‘Separate Universe’ technique, which can be employed to understand the response of small-scale galaxy physics to long-wavelength fluctuations. Furthermore, we extend our study beyond halos to galaxies by utilizing an advanced halo occupation distribution (HOD) model to approximate the luminous red galaxies (LRGs) population targeted by the Dark Energy Spectroscopic Instrument (DESI) survey DESI:2019jxc . We show that we are able to improve the uncertainty on $b_{\phi}$ and thus $f_{\rm NL}^{\rm loc}$ by more than 80% for all galaxy tracer samples considered in this study.
@@ II.1 AbacusSummit
We use the base resolution boxes of AbacusSummit, each of which contains 69123 particles in a $2\,h^{-1}{\rm Gpc}$ box, each with a mass of $M_{\rm part}=2.1\times 10^{9}\,h^{-1}{\rm M}_{\odot}$. While the AbacusSummit suite spans a wide range of cosmologies, here we focus on a single realization of the fiducial cosmology (Planck 2018: $\Omega_{b}h^{2}=0.02237$, $\Omega_{c}h^{2}=0.12$, $h=0.6736$, $2.083010^{9}A_{s}=$ 2.0830, $n_{s}=0.9649$, $w_{0}=-1$, $w_{a}=0$), AbacusSummit_c000_ph000 as well as two additional runs with a 2% higher and lower values of $\sigma_{8}$, AbacusSummit_c112_ph000 and AbacusSummit_c113_ph000. In particular, utilize the halo and particle catalogs (2022MNRAS.509..501H, ) as well as initial conditions outputs. For full details on all data products, see Ref. (2021MNRAS.508.4017M, ).
@@ II.2 Concentration definition
[…]
@@ II.3 Local environment
Another secondary parameter (beyond halo mass and linear bias) that affects both the halo and galaxy clustering is the environment of the halo. This effect has been studied extensively in the literature (2007MNRAS.378..641A, ; 2017A&A…598A.103P, ; 2018MNRAS.476.5442P, ; 2018MNRAS.473.2486S, ). A halo residing in a dense region is expected to contain more galaxies on average than a halo in an underdense region because halos in overdense regions experience more mergers, whereas those in underdense regions accrete mass more smoothly.
The response of galaxy formation in different halo environments in the presence of local-type PNG is of particular relevance to our study. To define halo environment, we adopt the same formalism as Ref. (2020MNRAS.493.5506H, ). Specifically, for each halo, we find all neighboring halos beyond its virial radius but within $r_{\rm max}=5\,h^{-1}{\rm Mpc}$ of the halo center. We sum the mass of all neighboring halos, denoted by $M_{\rm env}$, and define the environment parameter, $f_{\rm env}$:
@@ II.4 Separate Universe approach
We employ the separate universe technique McDonald:2001fe ; Sirko:2005uz ; Wagner:2014aka ; Lazeyras:2015lgp to measure the local-type PNG-induced bias, $b_{\phi}$, by invoking the equivalence between the response of galaxy formation to long-wavelength perturbations and the response of galaxy formation to changes in the background cosmology. The separate universe argument states that as long as the physics of galaxy formation acts on much smaller scales compared with the size of the long-wavelength perturbations, these long-wavelength perturbations effectively are equivalent to a universe with slightly modified background.Thus, the formation of galaxies at fixed cosmology in some region of space embedded in a long-wavelength fluctuation is equivalent to the formation of galaxies in a modified cosmology.
To access the quantity $b_{\phi}$, the modified cosmology needs to have a different amplitude of the primordial scalar power spectrum, $A_{s}$, or equivalently, the amplitude of linear fluctuations smoothed on a scale of $8\ {\rm Mpc}/h$, i.e., $\sigma_{8}$. Mathematically, the PNG-induced bias, $b_{\phi}$, is defined as
@@
where $\phi_{L}$ is to be treated as a constant locally and denotes the amplitude of the long-wavelength potential perturbation. Thus, we can evaluate $b_{\phi}$ in a separate universe with a different value of $A_{s}$ ($\sigma_{8}$), following the same procedure as (2024PhRvD.109j3530H, ), which we summarize below. In practice, we typically generate simulations with the same initial seed as the fiducial box, but with an input power spectrum file multiplied by $\left[1+\delta\sigma_{8}^{2}\right]$. In observations, making this measurement directly is extremely challenging, but recent works have suggested a possible path forward that directly extracts the PNG signal from the galaxy density field (2023PhRvD.107f1301G, ).
As in (2024PhRvD.109j3530H, ), we use the ‘Linear derivative’ AbacusSummit boxes, base_c112_ph000 and base_c113_ph000, which have the same initial seed as the fiducial simulation base_c000_ph000, but a different value of $\sigma_{8}$: namely, $\sigma_{8}^{\rm high}=1.02\times\sigma_{8}^{\rm fid}$ and $\sigma_{8}^{\rm low}=0.98\times\sigma_{8}^{\rm fid}$, respectively. We estimate $b_{\phi}$ in the separate universe approach as
@@
where $n_{h}^{\rm fid}(z)$, $n_{h}^{\rm high}(z)$ and $n_{h}^{\rm low}(z)$ is the mean number density of halos (galaxies) in the fiducial, high- and low-$\sigma_{8}$ simulations at redshift $z$.
@@ II.5 Hybrid Effective Field Theory
In this section, we summarize our approach for estimating the higher-order bias parameters, used to describe the connection between biased tracers such as galaxies and halos, and the underlying matter field in Lagrangian Perturbation Theory (LPT). For a review of perturbation theory, see Ref. 2002PhR…367….1B ; 2018PhR…733….1D ; 2020moco.book…..D . This standard version of LPT has been extended in the context of the Effective Theory of Large-Scale Structure Baumann:2010tm ; Carrasco:2012cv ; Senatore:2014eva ; 2009JCAP…08..020M in Lagrangian space 2014JCAP…05..022P ; 2015JCAP…09..014V ; Chen:2020fxs ; Chen:2020zjt .
In the Lagrangian picture, we work with infinitesimal fluid elements labeled by their initial (Lagrangian) positions $\bm{q}$. The dynamics is described by the displacement $\bm{\Psi}(\bm{q},\eta)$, generated by the gravitational potential and such that the Eulerian (comoving) positions $\bm{x}$ of the fluid element at some conformal time $\eta$ is $\bm{x}(\bm{q},\eta)=\bm{q}+\bm{\Psi}(\bm{q},\eta)$ 2015JCAP…09..014V ; 2014JCAP…05..022P .
More recently, a hybrid approach, Hybrid Effective Field Theory (HEFT), that leverages the strength of LPT as an analytical theory and the capability of $N$-body simulations to solve the full non-linear dynamics has been proposed and shown to perform very well in both simulations and observations (2020MNRAS.492.5754M, ; 2021MNRAS.505.1422K, ; 2021JCAP…09..020H, ; 2023MNRAS.524.2407Z, ; 2023MNRAS.520.3725P, ).
We adopt this framework here, and thus write the dependence of the galaxy field along its trajectory as an expansion to second order in the initial conditions adopting the Lagrangian framework 2009JCAP…08..020M ; 2015JCAP…11..007S ; 2014JCAP…08..056A ; 2016JCAP…12..007V :
@@
where $b_{1}$, $b_{2}$, $b_{s}$ and $b_{\nabla}$ are free bias parameters and $s_{L}^{2}=s_{ij}s^{ij}$ is the shear/tidal field, with $s_{ij}\equiv(\partial_{i}\partial_{i}/\partial^{2}-\delta_{ij}/3)\ \delta_{L}$. We have also included the lowest-order non-local term $\nabla^{2}\delta_{L}$. The functional can then be advected to the real-space (Eulerian) position $\bm{x}$ 2008PhRvD..77f3530M :
@@
To estimate the bias parameters, $b_{i}$, of our model, we minimize the difference between the model and the observable, $\delta_{\rm g}$. This is similar to the procedure adopted in (Schmittfull:2018yuk, ; 2022MNRAS.514.2198K, ; Baradaran:2024jlh, ), which we follow closely. In brief, the stochastic residual field $\epsilon(\mathbf{k})$, after removing deterministic contributions is given by:
@@
Minimizing the power spectrum of $\epsilon$, the bias parameters are given by:
@@
When fitting up to a maximum wavenumber, $k_{\rm max}$, this reduces to:
@@
Recently, in a similar vein to our work, a number of studies have adopted field-level estimates of the bias parameters associated with perturbation theory, to study various cosmological and astrophysical phenomena (see e.g., 2024arXiv241201888I, ; Shiferaw:2024ehr, ; 2024arXiv241201888I, ). Of particular relevance to our work are Refs. 2023JCAP…08..004S ; 2024PhRvD.110f3538I ; 2024arXiv241018039S , which apply similar field-level bias estimation techniques to the study of PNG.
@@ II.6 Halo Occupation Distribution
The HOD model that we adopt for both LRGs and QSOs is given by the formalism of Ref. (2005ApJ…633..791Z, ):
@@
where $M_{\mathrm{cut}}$ determines the minimum mass of a halo to host a central galaxy, $M_{1}$ sets the pivot scale of the power law of satellite occupation, $\sigma$ controls the steepness of the transition from 0 to 1 in the number of central galaxies, $\alpha$ is the power law index on the number of satellite galaxies, $\kappa M_{\mathrm{cut}}$ gives the minimum halo mass to host a satellite galaxy, $f_{\mathrm{ic}}$, which is a downsampling factor controlling the overall number density of the mock galaxies.
The velocity of the central galaxy is taken as the average velocity of the so-called “L2” subhalo (see 2022MNRAS.509..501H, ). For the satellite galaxies, the velocities are inherited from random halo particles. The analysis of Ref. (2023arXiv230606314Y, ) also includes velocity bias, which is necessary for modeling redshift-space clustering on small scales (e.g. 2015MNRAS.446..578G, ; 2022MNRAS.510.3301Y, ). For a detailed description of the galaxy population models, see Ref. (2024PhRvD.109j3530H, ).